Prepare for a data collection as follows:

Each of the team members is required to unwrap 3 packs of 45g of ‘Skittles original fruity candy’.



Count the number of candies with the same colour in each pack.
Let X be the number of red candies in each pack. Count the number of RED candies in each pack (e.g. 5, 4, 7: see TABLE 1).
Record the data obtained as in TABLE 1, TABLE 2 AND TABLE 3 below.

TABLE 1: Number of candies by colour and by team member
Colour
Member no.1
Member no. 2
Member no. 3
Member no. 4
Member no. 5
Total
green












yellow












red
5+4+7 = 16










orange












purple












Total












(3 Marks)
1
Expert's answer
2020-12-20T16:49:09-0500

TABLE 1: Number of candies by colour and by team member



1. Number of candy in each pack

Total number of candy opened by all members

= (128+127+127+127) = 509

Each member opens 3 packs. So total pack = 12

Hence candy per pack

"= \frac{509}{12} \\

= 42.4 \\

= 43"

2. Total no of candy by colour